### v2

# InstructSAM: A Training-Free Framework for Instruction-Oriented Remote Sensing Object Recognition

Abstract
Language-guided object recognition in remote sensing imagery is crucial for large-scale mapping and automated data annotation. However, existing open-vocabulary and visual grounding methods rely on explicit category cues, limiting their ability to handle complex or implicit queries that require advanced reasoning. To address this issue, we introduce a new suite of tasks, including Instruction-Oriented Object Counting, Detection, and Segmentation (InstructCDS), covering open-vocabulary, open-ended, and open-subclass scenarios. We further present EarthInstruct, the first InstructCDS benchmark for earth observation. It is constructed from two diverse remote sensing datasets with varying spatial resolutions and annotation rules across 20 categories, necessitating models to interpret dataset-specific instructions. Given the scarcity of semantically rich labeled data in remote sensing, we propose InstructSAM, a training-free framework for instruction-driven object recognition. InstructSAM leverages large vision-language models to interpret user instructions and estimate object counts, employs SAM2 for mask proposal, and formulates mask-label assignment as a binary integer programming problem. By integrating semantic similarity with counting constraints, InstructSAM efficiently assigns categories to predicted masks without relying on confidence thresholds. Experiments demonstrate that InstructSAM matches or surpasses specialized baselines across multiple tasks while maintaining near-constant inference time regardless of object count, reducing output tokens by 89% and overall runtime by over 32% compared to direct generation approaches. We believe the contributions of the proposed tasks, benchmark, and effective approach will advance future research in developing versatile object recognition systems. The code is available at https://VoyagerXvoyagerx.github.io/InstructSAM/.
1 Introduction
Object recognition in remote sensing imagery captures a vast array of objects and phenomena across diverse environments, providing rich information for supporting achieving the Sustainable Development Goals issued by the United Nations persello2022deep ; burke2021using , such as wildlife monitoring wu2023deep ; zheng2025beluga , poverty estimation ayush2021efficient ; meng2022count , and disaster response gupta2019creating . The recent advent of powerful vision-language models (VLMs), such as CLIP radford2021learning , has ushered in a new era of remote sensing oriented open-vocabulary object recognition algorithms (e.g. detection li2024toward ; pan2025locate and segmentation huang2025zori ). However, existing open-vocabulary approaches predominantly rely on explicit category cues, which restrict their capability to handle complex or implicit queries that demand advanced reasoning and contextual understanding. In other words, the rich diversity of visible objects in remote sensing images, due to the bird’s-eye view, predicates that any predefined, fixed list of categories is inevitably incomplete, limiting its practicality for open-ended analysis in the real world.
(a) Close-set
(b) Open-vocabulary
(c) Open-ended
(d) Open-subclass
Figure 1: Comparison of four task settings in object detection and segmentation. The sample is from DIOR Dataset li2020object (a) Close-set: Standard annotation with categories defined in the DIOR dataset. (b) Open-vocabulary: Instruction specifies which categories to detect and segment (e.g., “soccer field”, “football field”, “parking lot”). (c) Open-ended: Instruction requires detection and segmentation of all visible objects without category specification. (d) Open-subclass: Instruction targets all objects within a super-category (e.g., “sports fields”).
To address this issue, we expand the instruction-oriented object detection task pi2024ins and introduce a novel suite of tasks—Instruction-Oriented Object Counting, Detection, and Segmentation (InstructCDS) that encompasses open-vocabulary, open-ended, and open-subclass settings, as illustrated in Figure 1. The InstructCDS task entails a more flexible and scalable interpretation beyond fixed category sets and comprehends the complex users’ task requirements. We further present EarthInstruct, the first benchmark for InstructCDS in earth observation. The benchmark is constructed from two generic remote sensing object datasets covering 20 categories with different annotation rules and spatial resolution. EarthInstruct guides models to comprehend complex user instructions beyond the predefined three settings.
Recent advancements in VLMs have demonstrated impressive performance in object detection liu_grounding_2023 ; lin2024generative ; pi2024ins ; pan2025enhance, semantic segmentation yu2023convolutions ; cho2024cat ; zhang2025mirsam, visual grounding bai2025qwen2 ; chen2024expanding and reasoning-based segmentation  ren2024pixellm ; lai2024lisa within the natural image domain. However, transferring these methods to remote sensing imagery presents several challenges. First, direct inference leads to significant accuracy degradation due to the domain gap between natural and aerial images fu2024cross . Second, most existing remote sensing open-vocabulary detection li2024toward ; li2024exploiting ; zang2024zero ; li2025semantic and segmentation zermatten2025learning ; ye2025towards ; huang2025zori methods are trained on datasets with only a limited number of categories, restricting their generalization to diverse unseen categories. Third, conventional detectors li2022grounded ; cheng2024yolo ; minderer2023scaling rely on a threshold to filter predicted bounding boxes, which is not obtainable in zero-shot scenarios.
To tackle these challenges, we decompose the instruction-oriented object detection and segmentation tasks into several tractable steps and propose a framework without task-specific training, InstructSAM. First, a large vision-language model (LVLM) is employed to interpret user instructions and predict object categories and counts, with prompts systematically designed to maximize model’s capability. In parallel, SAM2 ravi2024sam is utilized to automatically generate mask proposals. Next, a CLIP model pre-trained on remote sensing images computes the semantic similarity between the predicted objects categories and mask proposals. We then formulate the object detection and segmentation as a mask-label matching problem, assigning predicted categories to mask proposals by integrating semantic similarity and global counting constraints. By inherently integrating three powerful foundation models, InstructSAM achieves superior performance across multiple tasks compared to both generic and remote sensing-specific VLMs trained on large-scale object recognition data. Notably, the inference time of InstructSAM remains nearly constant with respect to the number of predicted objects, reducing output tokens by 89% and overall runtime by 32% compared to directly generating bounding boxes using Qwen2.5-VL bai2025qwen2 in the open-ended setting. Our work paves the way for scalable, instruction-driven remote sensing object detection and segmentation, eliminating the need for costly pre-training or manual threshold tuning. Furthermore, the training-free paradigm allows InstructSAM to recognize objects in natural images when equipped with generic CLIP models.
In summary, our contributions are as follows:
  • We introduce the InstructCDS task, which challenges models to interpret user-provided natural language instructions and infer the number and locations of relevant objects.
  • We construct EarthInstruct to benchmark InstructCDS in earth observation, covering open-vocabulary, open-ended, and open-subclass settings and counting, detection, and segmentation tasks.
  • We develop InstructSAM, a training-free and confidence-free framework achieving near-constant inference time for the InstructCDS task.
  • Experiments on the benchmarks demonstrate that InstructSAM matches close-set models in object counting and surpasses generic and remote sensing-specific models in open-vocabulary and open-ended object recognition.
2 Related work
2.1 Instruction-oriented object detection and segmentation
Instruction-oriented methods
Instruction-oriented object detection (IOD), first introduced in pi2024ins , includes four instruction settings: category-specified (open-vocabulary wu2024towards ) detection, detection of all objects (open-ended lin2024generative ), detection within a super-category (we name it open-subclass), and detection to achieve certain goals. Ins-DetCLIP pi2024ins trains a detector to identify foreground objects and passes their features to a large language model (LLM) to generate categories based on user instructions. In addition to models specifically designed for the IOD task, Qwen2.5-VL bai2025qwen2 , trained on multi-task instruction data, also showcases the ability of dense object detection. However, both approaches require extensive task-specific training data, and their inference times increase substantially with the number of objects.
Open-vocabulary methods
Large-vocabulary object detection and segmentation datasets shao2019objects365 ; gupta2019lvis ; wang2023v3det and visual grounding datasets krishna2017visual ; hudson2019gqa enable various open-vocabulary learning approaches, including knowledge distillation gu2021open and region-text pre-training kamath2021mdetr ; minderer2022simple ; liu_grounding_2023 ; li2022grounded . Self-training with confidence threshold-filtered pseudo labels from image-text pairs (e.g., CC3M changpinyo2021conceptual , WebLI2B chen2022pali ) further boosts the performance minderer2023scaling ; liu_grounding_2023 ; zhang2022glipv2 ; yao2024detclipv3 . However, the quality of pseudo labels is highly sensitive to the chosen threshold minderer2023scaling , and these methods require predefined object categories, limiting their flexibility in diverse scenarios.
Open-ended methods
GenerateU lin2024generative first formulates the open-ended object detection (OED) problem. Concurrent works such as DetCLIPv3 yao2024detclipv3 , Florence-2 xiao2024florence , and DINO-X ren2024dino introduce generative frameworks that jointly predict object categories and bounding boxes using language models. However, constructing large-scale datasets with bounding box and caption pairs is resource-intensive. VL-SAM lin_training_free_2024 proposes a training-free approach via attention as prompts, but its iterative mask refinement and multi-prompt ensemble strategies are computationally expensive.
2.2 Instruction-oriented remote sensing object detection and segmentation
Recent advances in VLMs radford2021learning ; liu2023visual have also enabled open-vocabulary learning in the remote sensing domain. Diverse semantic tags from OpenStreetMap haklay2008openstreetmap and those generated by LVLMs drive the development of contrastive language-image pre-training in remote sensing images zhang2024rs5m ; wang2024skyscript . Following the generic open-vocabulary learning frameworks, methods for remote sensing open-vocabulary detection li2024toward ; wei2024ova ; pan2025locate ; huang2025openrsd and segmentation zermatten2025learning ; ye2025towards ; huang2025zori ; li2024segearth methods emerge. However, their human-annotated training data, being limited to a few dozen categories li2020object ; waqas2019isaid ; sun2022fair1m ; liu2025control, hinder generalization to out-of-distribution or zero-shot scenarios. Although some remote sensing LVLMs could support object recognition tasks like single-class object detection kuckreja2024geochat ; luo2024skysensegpt ; irvin2024teochat , visual grounding muhtar2024lhrs ; pang2025vhm , referring expression segmentation zhou2024geoground ; ou2025geopix , grounded conversation generation shabbir2025geopixel , and scene graph generation luo2024skysensegpt , they failed to follow complex reasoning instructions, such as open-vocabulary and open-subclass object detection. To annotate vast-vocabulary training data for remote sensing object detection, LAE-Label pan2025locate employs a generic LVLM chen2024far to predict categories for cropped mask proposals. However, this approach loses global context for accurate category classification.
In contrast, our InstructSAM adopts a confidence-free paradigm, requires no task-specific pre-training or finetuning, and maintains near-constant inference time regardless of object counts.
Figure 2: Annotation samples from NWPU-VHR-10 (a) and DIOR (b-d) illustrating rules that might differ from common sense. (a) Ships and vehicles are unannotated in low-resolution images (NWPU-VHR-10). (b) Expressway service areas separated by a road are treated as distinct instances (DIOR). (c, d) Airports are annotated only when fully visible (DIOR).
3 EarthInstruct, a benchmark for InstructCDS in remote sensing
3.1 Instruction setup
To enable practical applications in remote sensing such as large-scale mapping ayush2021efficient and image annotation, we define three distinct settings for InstructCDS:
1) Open-vocabulary: Counting, detection, and segmentation with user-specified categories.
2) Open-ended: Counting, detection, and segmentation of all objects without specifying categories.
3) Open-subclass: Counting, detection, and segmentation of objects within a super-category.
We construct EarthInstruct using NWPU-VHR-10 cheng2014multi and DIOR li2020object datasets, selected for their widespread use and diverse sensors, resolutions, and annotation rules. Critically, these dataset-specific annotation rules might deviate from common sense (e.g., excluding low-resolution vehicles) or exhibit semantic ambiguities (e.g., “bridge” vs. “overpass”), reflecting the original annotators’ specific goals (Figure 2). Consequently, simple instructions like “count vehicles” would fail to capture the nuances required by dataset conventions or user intent. EarthInstruct therefore necessitates models to interpret detailed instructions that clarify target definitions and handle dataset-specific rules (e.g., “do not count vehicles in images with a spatial resolution lower than 1m.”). To ensure fair evaluation aligned with dataset conventions and user requirements, prompts are designed accordingly, but image-specific prompts are prohibited to maintain scalability for large-area applications where prior content knowledge for each image is unavailable.
3.2 Evaluation metrics
Multi-class object counting
Standard counting metrics, such as Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE), used in benchmarks like FSC-147 ranjan2021learning and RSOC gao2020counting , inadequately capture nuanced multi-class evaluation. They cannot distinguish between over-counting and under-counting errors. Additionally, being unnormalized, they allow categories with larger counts to disproportionately skew the overall score when being averaged across classes.
To address these issues, we adopt precision, recall, and $\text{F}_{\text{1}}$-score, whereby offering normalized, per-class insights. We define per-image, per-class counting components as follows: let $C_{gt}$ denote the ground truth count and $C_{pred}$ the predicted count for a class in an image. Then, True Positives (TP) = $\min(C_{gt},C_{pred})$, False Positives (FP) = $\max(0,C_{pred}-C_{gt})$ for over-counting, and False Negatives (FN) = $\max(0,C_{gt}-C_{pred})$ for under-counting. These definitions enable standard calculation of precision, recall, and $\text{F}_{\text{1}}$-score, aggregated across images per class and then averaged for a final score.
Rethinking metrics for confidence-free detectors
Evaluating generative models like Florence-2 xiao2024florence or Qwen2.5-VL bai2025qwen2 , which output detections without confidence scores, poses a challenge for standard metrics. Average Precision (AP) everingham2010pascal relies on confidence scores to rank predictions and generate precision-recall curves. Without such ranking, standard AP is ill-defined. Furthermore, practical applications often involve filtering predictions with a fixed threshold, treating all remaining detections equally ayush2021efficient ; minderer2023scaling .
To resolve these issues and ensure fair comparison, we adopt confidence-free metrics: mean $\text{F}_{\text{1}}$-score ($\text{mF}_{\text{1}}$) and mean Average Precision with no confidence ($\text{mAP}_{\text{nc}}$li2025simple . $\text{mF}_{\text{1}}$ measures performance at a single operating point, suitable for fixed-threshold deployment. $\text{mAP}_{\text{nc}}$ adapts AP by assigning maximum confidence to all predictions. For confidence-free models, these metrics are computed directly (Results in Table 10). For conventional detectors (e.g., ren2016faster ; carion2020end ) that provide scores, when the confidence threshold is swept from 0 to 1 (step 0.02), the threshold maximizing $\text{mF}_{\text{1}}$ (using an IoU threshold of 0.5) across categories is selected, and the corresponding cusp score is reported.
Evaluating open-ended and open-subclass settings
In open-ended and open-subclass settings, LVLMs may generate category names (e.g., “car”) that differ textually from ground truth labels (e.g., “vehicle”). To handle this synonymy during evaluation, we adopt semantic similarity matching by following established protocols lin2024generative ; pi2024ins . Specifically, we encode generated categories and ground truth categories using the GeoRSCLIP zhang2024rs5m text encoder with the template ‘‘a satellite image of a {category}’’. A generated category name is deemed equivalent to a ground truth category if their embedding cosine similarity exceeds 0.95. This allows predicted objects associated with the generated name to be accurately evaluated against the matched ground truth category.
4 InstructSAM
Addressing the challenges of instruction-following, domain gaps, and threshold sensitivity in remote sensing object recognition, we propose a training-free framework named InstructSAM. It decomposes InstructCDS into three synergistic steps: instruction-based object counting using an LVLM, class-agnostic mask proposing via SAM2, and a novel counting-constrained mask-label matching procedure. This approach avoids the costly model training and threshold tuning, offering efficient and robust performance.
4.1 Instruction-oriented object counting with LVLM
As illustrated in Section 3, accurately interpreting user intent in remote sensing requires handling dataset-specific rules and semantic ambiguities, which the simple category prompts could fail to capture. We leverage state-of-the-art LVLMs (e.g., GPT-4o hurst2024gpt , Qwen2.5-VL bai2025qwen2 ) for this task. Inspired by he2024does , we utilize structured prompts in JSON format, which allows easy integration of dataset-specific Instructions alongside the core Task (detailed in Appendix C). Given an image $I$ and a detailed prompt $P$, the LVLM acts as a counter, outputting the target categories $\{cat_{j}\}$ and their corresponding counts $\{num_{j}\}$ present in the image: $\{cat_{j},num_{j}\}_{j=1}^{M}=\texttt{LVLM-Counter}(I,P)$.
4.2 Class-agnostic mask proposing
Concurrent with counting, SAM2 ravi2024sam is employed to generate high-quality, class-agnostic object masks for its strong generalization to remote sensing imagery wang2023samrs ; wu2023samgeo . Using its automated mask generation mode prompted by a regular point grid, we obtain a dense set of masks proposals $\{mask_{i}\}_{i=1}^{N}$. To enhance recall for small objects, mask generation is also applied to image crops (detailed in Appendix C.4).
4.3 Mask-label matching with counting constraints
A key innovation of InstructSAM is reframing object detection and segmentation as a constrained mask-label matching problem, by integrating the outputs from previous steps. Rather than using fragile confidence thresholds cheng2024yolo ; minderer2023scaling , we utilize global counts $\{num_{j}\}$ derived from the LVLM to constrain the assignment of semantic labels $\{cat_{j}\}$ to visual mask proposals $\{mask_{i}\}$.
Given $N$ mask proposals and $M$ target categories with counts, we compute a semantic similarity matrix $S\in\mathbb{R}^{N\times M}$, where $s_{ij}$ represents the cosine similarity between the CLIP zhang2024rs5m image embedding of a patch cropped around $mask_{i}$ (scaled by 1.2 for context) and the text embedding of $cat_{j}$ (using the template ‘‘a satellite image of a {category}’’). We then seek a binary assignment matrix $X\in\{0,1\}^{N\times M}$, where $x_{ij}=1$ assigns $mask_{i}$ to $cat_{j}$, by solving the Binary Integer Programming (BIP) problem:
$\min_{\mathbf{X}}\quad$ $\sum_{i=1}^{N}\sum_{j=1}^{M}(1-s_{ij})\cdot x_{ij}$ (1)
s.t. $\sum_{j=1}^{M}x_{ij}\leq 1,$ $\forall i\in\{1,\ldots,N\}$ (2)
$\sum_{i=1}^{N}x_{ij}=num_{j},$ $\forall j\in\{1,\ldots,M\},\quad\text{if }N\geq\sum_{j=1}^{M}num_{j}$ (3)
$\sum_{i=1}^{N}\sum_{j=1}^{M}x_{ij}=N,$ $\text{if }N<\sum_{j=1}^{M}num_{j}$ (4)
where constraint (2) ensures each mask is assigned to at most one category. Constraint (3) enforces that the number of assigned masks for each category matches the counts provided by the LVLM. Constraint (4) handles cases where the number of proposals is less than the total target count, ensuring all proposals are assigned.
As depicted in Figure 3, this BIP formulation elegantly fuses visual information, semantic information, and quantitative information. The visual information comes from the CLIP embeddings of mask proposals $\{v_{i}\}$, which contribute to $s_{ij}$. Semantic information is derived from the categories’ sentence embeddings $\{t_{j}\}$, also contributing to $s_{ij}$. Quantitative information from object counts $\{num_{j}\}$ serves as constraints in (3). The problem is efficiently solvable using open-source BIP solvers like PuLP mitchell2011pulp . The resulting non-zero entries in $X$ define the final set of recognized objects $\{(mask_{i},cat_{j})|x_{ij}=1\}$.
Figure 3: The InstructSAM framework. Given an input image and a structured counting prompt, the LVLM Counter extracts target categories $\{cat_{j}\}$ (semantic info) and their counts $\{num_{j}\}$ (quantitative info). Concurrently, the Mask Proposer generates mask proposals $\{mask_{i}\}$ (visual info). A CLIP model computes the similarity matrix $S$ between mask embeddings (from scaled crops) $\{v_{i}\}$ and category embeddings $\{t_{j}\}$. Finally, the Binary Integer Programming (BIP) Solver optimally assigns categories to masks by maximizing summed similarity, subject to the counting constraints, yielding the final recognition results.
5 Experiments
5.1 Implementation
We implement InstructSAM using GPT-4o-2024-11-20 hurst2024gpt (short for InstructSAM-GPT4o) or Qwen2.5-VL-7B bai2025qwen2 (short for InstructSAM-Qwen) as the LVLM counter, SAM2-hiera-large ravi2024sam for mask proposal, and GeoRSCLIP-ViT-L zhang2024rs5m for similarity computation. For the open-vocabulary setting, we follow previous works wei2024ova ; huang2025zori ; zang2024zero to split base and novel classes, and report $\text{mF}_{\text{1}}$, mean Insert-over-Union (mIoU), or $\text{mAP}_{\text{nc}}$. For open-subclass setting, we set two parent class “means of transport” and “sports fields”. We compare InstructSAM with a wide range of models, whose training data and abilities are listed in Table 6.
5.2 Results on EarthInstruct
Open-vocabulary setting
We report mean metrics across all classes for generic methods liu_grounding_2023 ; minderer2023scaling ; bai2025qwen2 and remote sensing open-vocabulary segmentation models ye2025towards ; li2024segearth trained on broader vocabularies in Table 1. Zero-shot performance on novel classes for models trained on base classes zang2024zero ; wei2024ova ; huang2025zori appears in Table 10. Models using novel class images or trained on entire detection datasets are evaluated on two additional datasets (Table 11).
Table 1 shows InstructSAM (esp. with GPT-4o) leading in all tasks with top counting metrics. On novel classes (Table 10), InstructSAM-Qwen achieves superior or competitive $\text{mAP}_{\text{nc}}$ against specialized models. This underscores InstructSAM’s training-free, counting-constrained matching advantage over traditional or fine-tuned approaches.
Table 1: Zero-shot performance comparison on EarthInstruct under open-vocabulary setting.
Method NWPU-VHR-10 DIOR
Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$ IoU Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$ IoU
Grounding DINO liu_grounding_2023 14.9 14.0 - - 10.7 6.0 - -
OWLv2 minderer2023scaling 39.4 27.2 - - 23.4 14.3 - -
Qwen2.5-VL bai2025qwen2 68.0 36.4 - - 52.0 27.8 - -
GSNet ye2025towards - - 1.3 6.4 - - 0.0 0.3
SegEarth-OV li2024segearth - - 3.5 12.1 - - 1.2 6.7
InstructSAM-Qwen 73.2 38.9 23.7 12.1 59.3 24.7 24.0 18.5
InstructSAM-GPT4o 83.0 41.8 26.1 14.8 79.9 29.1 28.1 20.2
Open-ended setting
Table 2: Performance comparison on EarthInstruct under open-ended setting.
Method NWPU-VHR-10 DIOR
Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$ Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$
Qwen2.5-VL bai2025qwen2 48.6 32.0 - 36.6 21.7 -
GPT-4o hurst2024gpt + OWL minderer2023scaling 32.6 24.0 - 30.6 21.0 -
SkySenseGPT luo2024skysensegpt 34.9 1.7 - 39.2 6.5 -
EarthDial soni2025earthdial 30.0 5.6 - 47.6 22.8 -
LAE-Label pan2025locate 46.2 27.3 24.1 23.3 11.5 10.4
GeoPixel shabbir2025geopixel 40.8 29.9 29.1 21.4 13.8 14.7
InstructSAM-Qwen 55.2 29.6 28.5 33.2 14.3 14.4
InstructSAM-GPT4o 57.4 31.3 29.9 47.9 22.1 21.8
Table 2 summarizes the results under open-ended setting. InstructSAM consistently achieves equivalent or higher $\text{F}_{\text{1}}$-scores than remote sensing-specific methods, including those trained on grounded description tasks luo2024skysensegpt ; shabbir2025geopixel . Notably, InstructSAM surpasses LAE-Label pan2025locate by leveraging a global view of the image to accurately predict object categories. While the absence of class-specific instructions in this setting limits further gains, InstructSAM still demonstrates robust performance (Figure 4).
Open-subclass setting
Table 3: Performance comparison on EarthInstruct under open-subclass setting. ‘S’ denotes the parent category “sports field”, and ‘T’ denotes “means of transport”.
Method NWPU-VHR-10 DIOR
Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$ Cnt $\text{F}_{\text{1}}$ Box $\text{F}_{\text{1}}$ Mask $\text{F}_{\text{1}}$
S T S T S T S T S T S T
Qwen2.5-VL bai2025qwen2 54.1 48.9 32.4 42.2 - - 52.2 51.8 34.0 39.2 - -
GPT-4o hurst2024gpt + OWL minderer2023scaling 40.3 68.0 19.8 65.9 - - 41.5 73.6 27.6 70.9 - -
InstructSAM-Qwen 50.9 55.3 33.5 41.9 33.0 39.7 40.4 67.4 22.2 49.3 22.6 49.5
InstructSAM-GPT4o 84.2 60.6 46.9 44.2 45.8 41.6 82.2 53.6 40.9 38.3 40.2 38.6
Table 3 shows that InstructSAM outperforms or matches Qwen2.5-VL across both parent classes. When prompted with categories identified by GPT-4o, OWLv2 performs strongly for “means of transport” but struggles with “sports fields”, likely due to the prevalence of transport-related categories in natural image datasets. These findings are consistent with the open-vocabulary results, where generic detectors, such as Grounding DINO and OWL, struggle with remote sensing categories except for airplane, vehicle, and ship.
(a) Open-vocabulary setting
(b) Open-ended setting
(c) Open-subclass setting
Figure 4: Qualitative comparison of object detection and segmentation methods on DIOR dataset. InstructSAM consistently outperforms other methods in generating accurate bounding boxes and segmentation masks across various settings. More qualitative results can be found in Appendix F.
Figure 5: Inference time as a function of bounding box count for InstructSAM-Qwen and baseline methods. Solid lines indicate linear regressions, and scatter points represent individual samples. The shaded regions for InstructSAM illustrate the time composition of different processing steps. Experiments are conducted on an RTX 4090 GPU.
Inference time analysis
Figure 5 compares inference times for open-ended methods using 7B LLMs. Mask proposal dominates the runtime of InstructSAM, whereas the PuLP optimizer solves the BIP problem in 0.07 seconds. InstructSAM maintains nearly constant inference time, while other methods scale linearly with the number of objects. Unlike representing bounding boxes as natural language tokens, InstructSAM reduces output tokens by 89% and total inference time by 32% compared to Qwen2.5-VL. This advantage will grow more pronounced as the model size scales up, highlighting the efficiency of our framework.
[…]
Method Add Instr NWPU vehicle NWPU all DIOR all
Faster-RCNN $\times$ 15/49/23 77/77/73 91/74/81
Qwen2.5VL $\times$ 11/75/19 80/72/72 73/53/56
Qwen2.5VL $\checkmark$ 28/70/41 82/71/73 83/51/59
GPT-4o $\times$ 11/80/20 68/79/67 87/65/72
GPT-4o $\checkmark$ 75/68/71 83/83/83 86/75/80
Table 5: Ablations on model scaling. MP recall is class-agnostic recall of mask proposals.
LVLM Counter Mask Proposal CLIP MP Recall Box $\text{F}_{\text{1}}$
GPT-4o SAM2-L DFN2B-L fang2024data 82.4 41.3
GPT-4o SAM2-L RemoteCLIP-L liu2024remoteclip 82.4 45.7
GPT-4o SAM2-L SkyCLIP-L wang2024skyscript 82.4 45.8
GPT-4o SAM2-L SkyCLIP-B wang2024skyscript 82.4 45.2
GPT-4o SAM2-S SkyCLIP-B wang2024skyscript 79.1 44.1
Qwen2.5VL SAM2-S SkyCLIP-B wang2024skyscript 79.1 40.6
5.3 Ablation studies
Prompt design
Table 4 reveals how additional instructions enhance object counting, particularly for categories with ambiguous or dataset-specific annotation rules. Initially, DIOR-trained Faster-RCNN and LVLM counters exhibit low vehicle precision on NWPU-VHR-10. Explicit annotation rules in instructions significantly boost vehicle precision for Qwen2.5-VL and GPT-4o, and also improve $\text{mF}_{\text{1}}$ by 3% and 8%, respectively, on DIOR. Contrary to zhang2024good , these results show that capable foundation models with precise, instruction-driven prompts truly empower LVLMs to match or exceed closed-set model performance.
Figure 6: Threshold sensitivity analysis in open-vocabulary setting. Left: Impact on mean metrics. Right: Category-specific $\text{F}_{\text{1}}$-scores. Dashed lines indicate optimal thresholds maximizing $\text{mF}_{\text{1}}$.
Model generalization and scaling
To assess InstructSAM’s generalization and scalability, we ablate LVLM counters, mask proposers, and CLIP models for the open-vocabulary detection (OVD) task on NWPU-VHR-10 (Table 5). InstructSAM consistently benefits from CLIP models fine-tuned on remote sensing data liu2024remoteclip ; wang2024skyscript over generic CLIP fang2024data , whereby yielding higher Box $\text{F}_{\text{1}}$-scores. Performance improves with larger model components, demonstrating the framework’s scalability. Notably, even with a smaller SAM2-S and SkyCLIP-B, InstructSAM coupled with Qwen2.5-VL (40.6 Box $\text{F}_{\text{1}}$) outperforms direct detection using Qwen2.5-VL alone (36.4 Box $\text{F}_{\text{1}}$), underscoring the efficacy of our approach.
Mask-label matching with counting constraints
Using fixed thresholds to filter CLIP predictions cheng2024yolo has inherent limitations. In Figure 6, the green and blue solid curves show InstructSAM without matching, where predictions are filtered by a similarity score (CLIP score) threshold before evaluation. Optimal thresholds vary widely across categories. As a result, a single global value performs poorly, and the results are highly threshold-sensitive, which is consistent with the findings in  li2025simple . In contrast, InstructSAM’s counting‑constrained matching removes this dependency by dynamically assigning predictions according to the estimated counts, yielding stronger performance in multi‑class and open‑world settings.
5.4 Error analysis of OVD task
Error identification reveals distinct error patterns across methods (Figure 7). OWLv2 primarily suffers from classification errors, while Qwen2.5-VL shows improved classification but struggles with missed detections. InstructSAM-GPT4o benefits from SAM2’s localization capabilities, though the background confusion persists due to GeoRSCLIP’s scene-focused training, which prioritizes broader contexts over individual objects.
(a) OWLv2
(b) Qwen2.5-VL
(c) InstructSAM
Figure 7: Error distribution of OVD task on NWPU-VHR-10 dataset. Pie charts show proportions of classification (Cls), localization (Loc), background (Bkg), and missed detection (Miss) errors across models.
6 Limitations
InstructSAM builds on pre-trained foundation models, and its performance inherits their capabilities and biases. For instance, SAM2 can miss the full extent of objects with intricate geometry (e.g., basketball court in Figure 4). CLIP models finetuned on scene-level remote sensing image–text pairs perform suboptimally when aligning object-level crops with semantic cues. More generally, foundation models such as GPT‑4o and SAM are trained primarily on optical imagery and struggle to count or segment objects in synthetic aperture radar (SAR) images.
Future work can mitigate these limitations by (i) incorporating foundation models trained on semantically diverse, multi-modal remote sensing data and (ii) leveraging advanced class-agnostic region proposers xie2025llama . Extending InstructSAM to semantic segmentation would further benefit the earth observation community. Although InstructSAM is only a first step, we aim to use it as an automated labeling engine to annotate at scale, enabling large, semantically rich object-recognition datasets.
7 Conclusions
In this paper, we introduce InstructCDS for instruction-driven object counting, detection, and segmentation, along with EarthInstruct, the first benchmark for this task in remote sensing domain. Our training-free InstructSAM framework integrates LVLMs, SAM2, and domain-specific CLIP to handle instruction-oriented scenarios with counting constrained mask-label matching. Experiments demonstrate that InstructSAM outperforms specialized baselines while maintaining near-constant inference time regardless of object counts. As the first approach extending instruction-oriented detection to the broader InstructCDS paradigm, InstructSAM will benefit from advancements in both remote sensing foundation models liu2025text2earth ; xiong2025geolangbind ; yan2023ringmo ; sahay2024aligning ; shan2025ros and general-purpose models team2025gemma ; zhu2025internvl3 ; luo2024mono , paving the way for more scalable, instruction-driven earth observation data analysis.
Acknowledgement
This work was supported by the National Earth Observation Data Center Research Project, National Natural Science Foundation of China (62506229), and Natural Science Foundation of Shanghai (25ZR1402268).
Appendix A Comparison with related work
Remote sensing imagery captures a vast, diverse range of contexts. Exemplified by the SkyScript dataset wang2024skyscript , curated from Google Earth and OpenStreetMap haklay2008openstreetmap , featuring 29,000 distinct semantic tags. Despite this richness, annotated training datasets for remote sensing object recognition typically span only a few dozen categories (Table 6). This lexical scarcity hinders effective alignment of semantic and visual information compared to generic models, which are typically trained on millions of human-annotated images and near-infinite vocabularies sourced from the Internet. Similarly, instruction-following datasets for remote sensing, such as FIT-RS luo2024skysensegpt and GeoPixelD shabbir2025geopixel , are also built upon manually annotated data, often inheriting these vocabulary-specific limitations.
Regarding model capabilities, remote sensing LVLMs typically fail to follow complex object recognition instructions as illustrated in Figure 8. Nevertheless, SkysenseGPT luo2024skysensegpt and GeoPixel shabbir2025geopixel , trained on scene graph generation or pixel-grounded conversation generation tasks, deliver comprehensive object detection or segmentation outputs, effectively serving a similar purpose of open-ended object recognition.
Table 6: Comparison of language-guided object recognition methods. We summarize the model, publication, training data, dense segmentation capability, and support for instruction-oriented object detection (open-vocabulary detection (OVD), open-ended detection (OED), and open-subclass detection (OSD)). Green indicates support, red indicates not supported.
Method Publication
Object Recognition
Training Data
Dense
Segmentation
OVD OED OSD
Remote sensing-specified methods
DesReg zang2024zero AAAI’24 DIOR
OVA-DETR wei2024ova arXiv’24 DIOR+DOTA+xView
CASTDet li2024toward ECCV’24 DIOR
LAE-DINO pan2025locate AAAI’25 LAE-1M
ZORI huang2025zori AAAI’25 NWPU/DIOR
GSNet ye2025towards AAAI’25 LandCover-40k
SegEarth-OV li2024segearth CVPR’25 Image-level Training
SkysenseGPT luo2024skysensegpt arXiv’24 FIT-RS
GeoPixel shabbir2025geopixel ICML’25 GeoPixelD
Generic methods
GroundingDINO liu_grounding_2023 ECCV’24 FiveODs+GoldG+Cap4M
OWLv2 minderer2023scaling NIPS’22 WebLI2B
Florence-2 xiao2024florence CVPR’24 FLA-5B
Qwen2.5-VL bai2025qwen2 arXiv’25 Various sources
InsDetCLIP pi2024ins ICLR’24 Object365
VL-SAM lin_training_free_2024 NeurIPS’24 Training-free
InstructSAM Ours Training-free
Figure 8: Examples illustrating that SkysenseGPT luo2024skysensegpt and TEOChat irvin2024teochat fail to produce meaningful responses for open-vocabulary, open-ended, and open-subclass prompts. Their responses either lack category name outputs or exhibit looped generation.
Appendix B Benchmark details
The details of the datasets comprising EarthInstruct are as follows:
  • NWPU-VHR-10 cheng2014multi is a 10-class dataset for very-high-resolution (VHR) remote sensing object detection. It originally comprises 650 positive images (each containing at least one target object) and 150 negative images (containing no target objects). For constructing EarthInstruct, we utilize the 650 positive images. Instance-level mask annotations were subsequently provided by su2019object .
  • DIOR li2020object is a large-scale benchmark for object detection in optical remote sensing images, encompassing 20 object categories. The dataset is split into 5,862 training images, 5,863 validation images, and 11,725 test images. For the EarthInstruct benchmark, we utilize the validation set of DIOR. Mask annotations are provided by the SAMRS wang2023samrs dataset.
Table 7 provides a summary of these datasets, including their spectral bands, spatial resolution ranges, and the specific category splits within EarthInstruct.
Table 7: Details of the datasets used in EarthInstruct, including spectral bands, resolution, and category splits for base, novel, and open-subclass (sports field, means of transport) settings. “Res” denotes spatial resolution. “NWPU” stands for NWPU-VHR-10.
Dataset Band Res Base Class Novel Class Sports Field Means of Transports
NWPU RGB/ Color Infrared 0.08-2m airplane, storage_tank, baseball_field, tennis_court, track_field, bridge, vehicle ship, basketball_ court, harbor baseball_field, tennis_court, basketball_court, track_field airplane, ship, vehicle
DIOR RGB 0.3-30m airplane, baseball field, bridge, chimney, expressway service area, expressway toll station, dam, golf field, harbor, overpass, ship, stadium, storage tank, tennis court, train station, vehicle airport, basketball court, ground track field, windmill baseball field, basketball court, golf field, ground track field, tennis court, stadium airplane, ship, vehicle
Table 8 compares EarthInstruct with other language-guided object recognition datasets and benchmarks. Most current benchmarks pi2024ins ; liu2024rrsisD ; li2024vrsbench generate prompts using LLMs or templates based on specific objects in the image, resulting in few negative samples. In addition, referring segmentation liu2024rrsisD and visual grounding li2024vrsbench benchmarks typically support only single-class or single-object queries, restricting efficiency in large-scale applications.
EarthInstruct addresses these issues by supporting open-vocabulary, open-ended, and open-subclass settings, and uses dataset-specific prompts that encode user intent without relying on image content priors. This enables more comprehensive and efficient evaluation for instruction-driven object recognition.
Table 8: Comparison between EarthInstruct and other language-guided object recognition benchmark datasets. “OV”, “OE”, “OS” denote open-vocabulary, open-ended, and open-subclass, respectively. Goal-Oriented denotes instructions to detect objects that achieve certain goals. Content Prior Free means the benchmark does not require image content prior for prompt construction.
Dataset Publication OV Instructions OE Instructions OS Instructions Goal- Oriented Instructions Content Prior Free Class Num
IOD-Bench pi2024ins ICLR’24
RRSIS-D liu2024rrsisD CVPR’24 20
VRSBench li2024vrsbench NeurIPS’24 26
FIT-RS (OD) luo2024skysensegpt arXiv’24 48
FIT-RS (SGG) luo2024skysensegpt arXiv’24 48
GeoPixelID shabbir2025geopixel ICML’25 15
EarthReason li2025segearth arXiv’25 28
EarthInstruct Ours 20
Appendix C Implementation details
This section provides further details on the implementation of our InstructSAM framework and other baselines. A critical component of leveraging LVLMs for the InstructCDS tasks is the design of effective prompts. We employ structured prompts in JSON format to systematically guide the LVLM counter. This structured approach allows for clear delineation of the model’s Persona, the specific Task to be performed, detailed Instructions for execution, the desired Output format, and illustrative Examples. The Instructions field enables the incorporation of dataset-specific annotation rules and disambiguation of category definitions, which are crucial for accurate interpretation the user intension, as discussed in Section 3.1. The subsequent subsections detail the specific prompts used for each setting within the EarthInstruct benchmark.
Here we further clarify how the categories that might have bewildered the model. For example, an “overpass” is defined as “a bridge, road, railway or similar structure that is over another road or railway [^1] https://en.wikipedia.org/wiki/Overpass”, while a “bridge” is “a structure built to span a physical obstacle (such as a body of water, valley, road, or railway) without blocking the path underneath [^2] https://en.wikipedia.org/wiki/Bridge.” It is possible that the LVLM counter will recognize overpasses as bridges since overpass is a subset of bridge. Adding clarification such as “bridge in this dataset refers to a structure spanning a body of water” to the Instructions field will guild the LVLM to count accurately.
C.1 Prompts in open-vocabulary setting
  • Prompt for NWPU-VHR-10 dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing aerial images.",
    "Task": "Given an input satellite imagery, count the number of objects from specific categories. Provide the results in JSON format where the keys are the category names and the values are the corresponding counts.",
    "Instructions": [
    "The 10 categories in the dataset are: [‘airplane’, ship’, storage_tank’, baseball_field’, tennis_court’, basketball_court’, track_field’, harbor’, bridge’, vehicle’]",
    "The spatial resolution of the imagery in the dataset ranges from 0.08 m to 2 m.",
    "Do not count ships or vehicles that are hard to annotate in the relatively low-resolution images as they are not annotated due to the small size.",
    "Harbor is defined as a pier to dock ships. If multiple harbors are visible in the image, count each distinct pier separately."
    "Output format": "{\"category1\": count1, \"category2\": count2, ... }",
    "Examples": [
    { "airplane": 2, "ship": 0, "storage_tank": 3, "baseball_field": 1, "tennis_court": 0, "basketball_court": 0, "track_field": 0, "harbor": 6, "bridge": 0, "vehicle": 0 },
    { "airplane": 5, "ship": 2, "storage_tank": 0, "baseball_field": 0, "tennis_court": 1, "basketball_court": 0, "track_field": 0, "harbor": 0, "bridge": 1, "vehicle": 10 }
  • Prompt for DIOR dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, count the number of objects from specific categories. Provide the results in JSON format where the keys are the category names and the values are the corresponding counts.",
    "Instructions": [
    "The 20 categories in the dataset are: [‘airplane’, airport’, baseball field’, basketball court’, bridge’, chimney’, expressway service area’, expressway toll station’, dam’, golf field’, ground track field’, harbor’, overpass’, ship’, stadium’, storage tank’, tennis court’, train station’, vehicle’, windmill’]",
    "The spatial resolution of the images is 0.3m-30m.",
    "Airport is a large area of land where aircraft can take off and land. It includes runways and other facilities. Do not count airport if the it is not compeletely visible in the image.",
    "Harbor is defined as a pier to dock ships. If multiple harbors are visible in the image, count each distinct pier separately.",
    "Expressway toll station is a toll booth at the entrance of the expressway and spans the road.",
    "Expressway service area is a rest area along an expressway. If it exsits on the both sides of the expressway, count them separately.",
    "Overpass is a road crossing over another road. Bridge is a road spanning a river. Distinguish them carefully.",
    "If the overpass or bridge is composed of parallel, separate sections (for example, different lanes or directions of traffic), each section should be counted individually.",
    "Count every ship and vehicle carefully, even the resolution is low and the objects are small and dense.",
    "If none of the objects among the categories is visible, output a JSON object with all categories set to 0"
    "Output format": "{ "category1": count1, "category2": count2, ... }",
    "Examples": [
    "{ "airplane": 2, "airport": 0, "baseball field": 0, "basketball court": 0, "bridge": 1, ... }",
    "{ "airplane": 0, "airport": 0, "baseball field": 2, "basketball court": 6, "bridge": 0, ... }"
C.2 Prompts in open-ended setting
  • Prompt for NWPU-VHR-10 dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images.",
    "Task": "Given an input satellite imagery, count the number of all the visible remote sensing objects. Provide the results in JSON format where the keys are the category names and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the imagery in the dataset ranges from 0.08 m to 2 m.",
    "Do not count ships or vehicles that are too samll and are hard to annotate in the relatively low-resolution images.",
    "Only count objects that are clearly visible in the imagery. If a category is not visible, do not include it in the output."
    "Output format": "{ "category1": count1, "category2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
  • Prompt for DIOR dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images.",
    "Task": "Given an input satellite imagery, count the number of all the visible remote sensing objects or scenes. Provide the results in JSON format where the keys are the category names and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the imagery in the dataset ranges from 0.3 m to 30 m.",
    "If the resolution is too limited or the scene is too dense to accurately count certain objects, exclude those objects from the results.",
    "Only count objects that are clearly visible in the imagery."
    "Output format": "{ "category1": count1, "category2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
C.3 Prompts in open-subclass setting
  • Prompt for counting “sports field” in NWPU-VHR-10 dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, count the number of objects that belong to the parent category **sports field**. Provide the results in JSON format where the keys are the names of the subcategories and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the images is 0.08m-2m.",
    "Do not count objects that are hard to annotate in the relatively low-resolution images as they are not annotated due to the small size.",
    "If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
    "Output format": "{ "subcategory1": count1, "subcategory2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
  • Prompt for counting “means of transport” in NWPU-VHR-10 dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, count the number of objects that belong to the parent category **means of transport**. Provide the results in JSON format where the keys are the names of the subcategories and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the imagery in the dataset ranges from 0.08 m to 2 m.",
    "Do not count boats or land vehicles that are hard to annotate in the relatively low-resolution images as they are not annotated due to the small size.",
    "If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
    "Output format": "{ "subcategory1": count1, "subcategory2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
  • Prompt for counting “sports field” in DIOR dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, count the number of visible objects that belong to the parent category **sports field**. Provide the results in JSON format where the keys are the names of the subcategories and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the images is 0.3m-30m.",
    "Return only the categories and counts that meet the visibility and resolution criteria. If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}".
    "Output format": "{ "subcategory1": count1, "subcategory2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
  • Prompt for counting “means of transport” in DIOR dataset:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, count the number of visible objects that belong to the parent category **means of transport**. Provide the results in JSON format where the keys are the names of the subcategories and the values are the corresponding counts.",
    "Instructions": [
    "The spatial resolution of the images is 0.3m-30m.",
    "Return only the categories and counts that meet the visibility and resolution criteria. If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
    "Output format": "{ "subcategory1": count1, "subcategory2": count2, ... }",
    "Answer": [
    "Ensure the category names are in singular form",
    "Provide the counts as integers."
C.4 Hyperparameters of InstructSAM
Mask proposer
As described in Section 4.2, class-agnostic mask proposals are generated using SAM2-hiera-large ravi2024sam in its automatic mask generation mode. The model was configured with the parameters detailed in Table 9 across both NWPU-VHR-10 and DIOR. These settings balance the proposal quality and computational efficiency, ensuring a dense set of high-quality proposals, especially for small objects.
Table 9: Hyperparameters for SAM2.
Parameter Value
pred_iou_thresh 0.75
stability_score_thresh 0.75
points_per_side 24
crop_n_layers 1
box_nms_thresh 0.5
LVLM counter
For Qwen2.5-VL and GPT-4o, the temperature is set at 0.01 and top_p is set at 1 to reduce randomness.
C.5 Implementations of baseline methods
To ensure rigorous and fair comparisons against InstructSAM, baseline methods are implemented or adapted as detailed below. Unless otherwise specified, publicly available pre-trained models and official codebases are utilized, adhering to their recommended configurations.
Open-Vocabulary semantic segmentation methods
For SegEarth-OV li2024segearth and GSNet ye2025towards , we follow their prescribed inference procedures. This involves providing the target categories along with an explicit “background” category as input. To derive instance-level proposals from the pixel-level segmentation maps produced by these models, we apply a standard connected component labeling algorithm to the per-class binary masks. This allows evaluation using instance-based metrics (Mask $\text{F}_{\text{1}}$).
Open-ended methods
  • SkysenseGPT luo2024skysensegpt : We utilize the prompt “[grounding]Analyze and describe every detail you can identify in the image,” which is a representative of its training data, and set the maximum output token limit to 5000.
  • GeoPixel shabbir2025geopixel : Following the illustrative examples provided in their publication, we employ the prompt “Can you give a thorough description of this image, including interleaved segmentation masks to highlight key objects?”
  • LAE-Label pan2025locate : For mask proposing, we use SAM2-hiera-large with the same hyperparameters as InstructSAM (see Appendix C.4). For label generation, we replace LAE-Label’s original InterVL2-8B and reasoning prompt with the newer Qwen2.5-VL-7B, guided by a structured prompt for open-ended category identification without explicit reasoning. This change of prompt improves $\text{mF}_{\text{1}}$ by 4.3% on NWPU-VHR-10 and reduces inference time by 92%. The prompt we use is as follows:
    "Persona": "You are an advanced AI model capable of understanding and analyzing remote sensing images."
    "Task": "Given an input satellite imagery, identify the most likely object class visible in the image. Provide the result as a single class name."
    "Instructions": [
    "Consider this is a region of interest cropped from a larger remote sensing image.",
    "Focus on identifying the main object class visible in the cropped region.",
    "Be specific with your answer, using a single class name."],
    "Output format": "\"class_name\"",
    "Answer": ["Provide only the class name in quotes, without additional explanations. If it is not recognized, output \"Unrecognized\" "]
Qwen2.5-VL
Given Qwen2.5-VL’s bai2025qwen2 extensive pre-training on object detection tasks, we request Qwen2.5-VL to generate a prompt template to perform IOD, and specify the output format following the examples in their paper. Taking DIOR dataset as an example, the prompts for each task are as follows.
  • Open-vocabulary detection:
    **Prompt:**
    I need assistance with performing an remote sensing object detection task using the provided category names: [‘airplane’, airport’, baseball field’, basketball court’, bridge’, chimney’, expressway service area’, expressway toll station’, dam’, golf field’, ground track field’, harbor’, overpass’, ship’, stadium’, storage tank’, tennis court’, train station’, vehicle’, windmill’].
    Please provide the results in JSON format, where each object is represented as follows:
    - Each object should include a label (category name) and its corresponding bounding box coordinates (in the format [x1, y1, x2, y2]).
    **Example Output:**
    {"label": "category_name", "bbox_2d": [x1, y1, x2, y2]},
    {"label": "category_name", "bbox_2d": [x1, y1, x2, y2]},
  • Open-ended detection:
    **Prompt:**
    Please detect all the visible objects in the satellite image and provide the results in JSON format, where each object is represented as follows:
    - Each object should include a label (category name) in singular form and its corresponding bounding box coordinates (in the format [x1, y1, x2, y2]).
    **Example Output:**
    {"label": "category1", "bbox_2d": [x1, y1, x2, y2]},
    {"label": "category2", "bbox_2d": [x1, y1, x2, y2]},
  • Open-subclass detection (e.g., for “sports field”):
    **Prompt:**
    Please detect all the visible objects in the satellite image that belong to the parent category **sports field** and provide the results in JSON format, where each object is represented as follows:
    - Each object should include a label (subcategory name) in singular form and its corresponding bounding box coordinates (in the format [x1, y1, x2, y2]).
    - If none of the objects belong to the parent category is visible, output a empty empty list like [].
    **Example Output:**
    {"label": "subcategory1", "bbox_2d": [x1, y1, x2, y2]},
    {"label": "subcategory2", "bbox_2d": [x1, y1, x2, y2]},
While incorporating detailed dataset-specific instructions into prompts enhances Qwen2.5-VL’s performance for object counting, such elaborations does not benefit direct detection. In fact, such instructions reduce $\text{mF}_{\text{1}}$ by 1% on NWPU-VHR-10 and have little effect on DIOR. The main paper reports open-vocabulary detection results without these instructions.
Appendix D Additional experiments
D.1 Comparison with zero-shot remote sensing object recognition methods on EarthInstruct
Direct comparison with some zero-shot remote sensing detection and segmentation methods (e.g., DesReg zang2024zero , OVA-DETR wei2024ova , ZoRI huang2025zori ) is difficult due to differing dataset splits, evaluation metrics, and limited reproducibility. Table 10 shows their reported mAP scores alongside InstructSAM’s $\text{mAP}_{\text{nc}}$ on novel classes. Notably, InstructSAM-Qwen matches or outperforms these methods on their respective benchmarks (e.g., ZoRI on NWPU-VHR-10, OVA-DETR on DIOR). While these baselines use confidence-based AP, InstructSAM’s confidence-free APnc results underscore its robustness.
Table 10: Comparison with zero-shot remote sensing object detection and segmentation methods on novel classes. ‘-’ means data missing due to limited reproducibility. $\emptyset$ indicates model lacking the segmentation ability. APnc is reported for InstructSAM, while values in gray for other methods are AP reported in their original papers.
Method NWPU-VHR-10 DIOR val DIOR test
Box AP Mask AP Box AP Mask AP Box AP Mask AP
DesReg zang2024zero - $\emptyset$ - $\emptyset$ 7.9 $\emptyset$
OVA-DETR wei2024ova - $\emptyset$ 7.1 $\emptyset$ - $\emptyset$
ZoRI huang2025zori - 12.3 - - - 8.5
InstructSAM-Qwen 24.6 24.1 7.6 6.3 4.9 4.3
InstructSAM-GPT4o 26.8 26.5 11.2 8.9 6.6 5.4
D.2 Comparison with open-vocabulary detection methods on out-of-distribution datasets
To compare with open-vocabulary methods that include DIOR into training, we compare the zero-shot performance on two out-of-distribution (OOD) datasets:
  • xBD gupta2019creating is a large-scale building damage assessment dataset with spatial resolution below 0.8m. The test set includes 933 pre- and post-disaster image pairs with instance-level building masks. We use pre-disaster images to evaluate object counting and detection with the single category “building”.
  • Aerial Maritime Drone Large li2022elevater is a drone-based object detection dataset with 74 aerial maritime images and 1,151 bounding boxes. Categories include “docks”, “boats”, “lifts”, “jetskis”, and “cars”. The entire dataset is used for zero-shot evaluation.
Table 11 shows that LAE-DINO, trained on LAE-1M (including “building”), achieves the highest detection $\text{F}_{\text{1}}$-core (50.6) on xBD. On Aerial Maritime Drone, however, LAE-DINO detects only boats, even struggling with docks and cars despite semantic similarity to harbor and vehicle categories in training. Moreover, the confidence-based detectors (OWL, CASTDet, LAE-DINO) require dataset-specific thresholds to optimize zero-shot results, highlighting InstructSAM’s advantage in eliminating threshold tuning.
Table 11: Comparison with open-vocabulary detection methods on xBD and Aerial Maritime Drone Large datasets.
Method xBD Aerial Maritime Drone Large
Best Thr Cnt $\text{mF}_{\text{1}}$ Det $\text{mF}_{\text{1}}$ Best Thr Cnt $\text{mF}_{\text{1}}$ Det $\text{mF}_{\text{1}}$
OWL minderer2023scaling 0.02 53.0 31.5 0.26 28.3 21.5
CASTDet li2024toward 0.00 0.0 0.0 0.40 13.3 10.8
LAE-DINO pan2025locate 0.08 79.6 50.6 0.30 13.0 12.5
InstructSAM-GPT4o - 65.9 37.2 - 44.5 22.5
D.3 Selection of prompt format
We conduct five independent experiments evaluating Qwen2.5-VL and GPT-4o’s open-vocabulary counting ability on NWPU-VHR-10 using Markdown and JSON prompts. Table 12 shows the mean accuracies and standard deviations. Qwen2.5-VL exhibits only a 1.2% difference in mean accuracy between Markdown and JSON, with zero standard deviation for both. GPT-4o shows slight variability: 82.38% ± 0.94 (Markdown) and 82.68% ± 0.39 (JSON). These results indicate that these LVLM counters maintain relatively stable performance across different prompt formats.
Table 12: Mean accuracy and 1-sigma standard deviation for different prompt formats and models.
Model Prompt Format Mean Accuracy (%) 1-Sigma (Stand Deviation)
Qwen2.5-VL Markdown 72.0 0.00
Qwen2.5-VL JSON 73.2 0.00
GPT-4o Markdown 82.4 0.94
GPT-4o JSON 82.7 0.39
Appendix E Broader impacts
InstructSAM demonstrates a strong ability to recognize objects in remote sensing imagery with versatile instructions. It accelerates large-scale mapping applications, further supporting public agencies and humanitarian organizations in critical areas such as poverty mapping, disaster response. Its training-free paradigm and near-constant inference speed lower computational costs, decreasing carbon emissions and broadening access to remote sensing applications in resource-limited settings.
InstructSAM’s capabilities also raise privacy concerns as the spatial resolution of satellite imagery is growing higher. Hallucinations or misclassifications, stemming from biases in pre-trained models may produce unreliable outputs, require user verification to ensure accuracy.
Appendix F Qualitative results
This section presents qualitative results of InstructSAM across open-vocabulary (Figure 9), open-ended (Figure 10), and open-subclass settings (Figure 11). Additionally, we showcase its versatility in handling instructions to achieve certain goals in Figure 12 and its generalization to natural images in Figure 13.
F.1 Qualitative results in open-vocabulary setting
Figure 9: Qualitative results in open-vocabulary setting. While OWLv2 struggles to distinguish remote sensing objects beyond vehicles, and SegEarth-OV fails to separate foreground objects from the background, InstructSAM demonstrates superior performance in segmenting remote sensing objects.
F.2 Qualitative results in open-ended setting
Figure 10: Qualitative results in open-ended setting. Unlike GeoPixel and SkysenseGPT, which fail to detect classes outside their training set, InstructSAM demonstrates its ability to recognize diverse objects (e.g., pavilion in (a), tree in (c)) and provides more accurate bounding boxes and less fragmented masks.
F.3 Qualitative results in open-subclass setting
(a) Detecting the subclass of “means of transport”
(b) Detecting the subclass of “sports field”
(c) Detecting the subclass of “sports field”
Figure 11: Qualitative results in open-subclass setting. InstructSAM effectively identifies objects within parent categories. In contrast, Qwen2.5-VL struggles with dense objects, and OWLv2 faces challenges in classifying sports fields from a bird’s-eye view.
F.4 Qualitative results on other instructions
Figure 12: Qualitative results on following versatile instructions. Samples are from EarthReason li2025segearth dataset. InstructSAM successfully recognizes objects based on implicit cues.
F.5 Qualitative results in natural images
Figure 13: Qualitative results on natural images. When equipped with generic CLIP models (DFN2B-CLIP fang2024data ), InstructSAM effectively recognizes objects in natural images.